OpenStudy (anonymous):

log(sqroot of) x+y / 10 EXPAND.

OpenStudy (anonymous):

Is ((x + y)/10) ALL under the square root sign, or just the (x+y) part? There are rules for logarithms. Look them up. eg. log x^2 = 2 log x. Square root is the same as an exponent of 1/2. eg. log 10 = 1 because 10^1 = 10 Also you might need the fact that log(a/b) = log(a) - log(b). If the 10 is part of the square root: (1/2)[log(x + y) - log10]

OpenStudy (anonymous):

Then you'd simplify to (1/2)(log(x + y) - 1). Or you could write it as (1/2)(log(x+y)) - 1/2 BUT if 10 is NOT part of the square root, then I'd start with: log (sqrt(x+y)) - log 10. So (1/2)(log(x+y)) - 1 would be the answer. The two answers are only a bit different.

OpenStudy (anonymous):

it's whole sqroot of x+y

OpenStudy (anonymous):

So if it's [log (x+y)] /10 you do this: log (sqrt(x+y)) - log 10. So (1/2)(log(x+y)) - 1 would be the answer.